### v1

# Growth, discovery and characterization of single crystalline Eu0.8Pt6Al15.8

Abstract
We report the discovery of a ternary compound, Eu0.8Pt6Al15.8. We determine its chemical and structural characteristics based on energy-dispersive X-ray spectroscopy as well as both powder and single-crystal X-ray diffraction, demonstrating that it crystallizes in a hexagonal structure type EuPt6Al16 with no reported structural analog. The electronic and magnetic properties are characterized by temperature- and field-dependent magnetization, and temperature-dependent resistance measurements, revealing that the Eu2+ magnetic moments order antiferromagnetically below 2.8 K.
keywords:
Crystal , Europium , Magnetism , Valence
journal: Journal of Solid State Chemistry
[…]
organization=Department of Physics and Astronomy, Iowa State University,addressline=, city=Ames, postcode=50011, state=IA, country=USA
[…]
organization=Ames National Laboratory, Iowa State University,addressline=, city=Ames, postcode=50011, state=IA, country=USA
[…]
organization=Universität des Saarlandes, Anorganische Festkörperchemie,addressline=Campus C4 1, city=Saarbrücken, postcode=66123, country=Germany
[…]
organization=Institut für Anorganische und Analytische Chemie, Universität Münster,addressline=Corrensstrasse 20, city=Münster, postcode=48149, country=Germany
[…]
organization=Department of Physics and Astronomy, Iowa State University,addressline=, city=Ames, postcode=50011, state=IA, country=USA
[…]
organization=Ames National Laboratory, Iowa State University,addressline=, city=Ames, postcode=50011, state=IA, country=USA
[…]
organization=Department of Physics and Astronomy, Iowa State University,addressline=, city=Ames, postcode=50011, state=IA, country=USA
[…]
organization=Ames National Laboratory, Iowa State University,addressline=, city=Ames, postcode=50011, state=IA, country=USA
1 Introduction
An extensive number of compounds with rare-earth, transition metals and aluminum with diverse structures have been found [1], many of them remain underexplored. In particular, a plethora of rare-earth ($R$) aluminum platinides have been discovered, such as RRPtAl [2], Ce3Pt4Al6 [3], RRPt6Al3 [4], $R_{2}$Pt9Al16 [5], and RRPtAl2 [6], and a broader class of RRPt5-xAlx which includes RRPt3Al2 as well as RRPt4Al [7]. Additionally, there have been reports of compounds built of disordered $R$-Al layers and a substantial amount of vacancies in the $R$ sites, such as $R_{1.33}$Pt3Al8 and $R_{0.67}$Pt2Al5, which suggests that a larger cell may be needed to describe these compounds [8]. Both of these have related structures made of the same $R$-Al layers but in an ordered arrangement forming the superstructures $R_{4}$Pt9Al24 [9] and $R_{2}$Pt6Al15 [10], respectively.
Recently, a Eu2Pt6Al15 phase was reported to exhibit a valence transition from Eu2+ to Eu3+ upon cooling below 45 K [11], according to Mössbauer spectroscopy, resistance and magnetic susceptibility measurements. This phase was obtained in polycrystalline form by arc-melting Eu, Pt and Al together, followed by a thermal treatment. As part of an attempt of obtaining that phase in single crystalline form by solution growth [12], we discovered a different ternary: Eu1-xPt6Al16-y, with $x\sim 0.2$ and $y\sim 0.2$. Since it adopts a structure that has not been reported in any other compound, in this paper we report the details for obtaining single crystals of this phase, as well as its basic structural, chemical, and magnetic characterization.
2 Experimental details
Single crystals of Eu0.8Pt6Al15.8, were obtained by the high-temperature solution growth method [13] out of a ternary melt with an initial nominal composition of Eu2Pt10Al88. The starting materials were elemental Pt powder (Ames National Laboratory, 99.9+% purity), pieces of Al rod (Alfa Aesar, 99.999% purity) and Eu pieces (Materials Preparation Center - Ames National Laboratory, 99.99+% purity).
There is a large exothermic reaction between Al and Pt/Pd upon melting from separate elements into a single-phase liquid. To prevent this exothermic reaction from occurring inside the alumina crucible in which the crystals grow, and potentially cracking the former, a step similar to that reported for the growth of the Al-Pd-Mn single grain quasicrystal [14] was followed. This consisted of pre-melting the Al and Pt using an arc furnace. Pt powder was first pressed into a pellet, arc-melted, weighed and then arc-melted in contact with Al pieces, to form a single button with composition of Pt10Al88. The arc-melted button was placed in a 2 ml alumina fritted Canfield Crucible Set [15, 16] together with the corresponding amount of Eu to achieve the desired initial composition. The crucible with the elements was sealed in a fused silica tube with a partial pressure of $\sim 1/6$ atm of Ar and placed inside a box furnace. It was heated to 1180 C, kept at that temperature for 5 hours, and slowly cooled down to 900 C over the course of 96 hours. At 900 C the excess solution was decanted with the aid of a centrifuge [13]. Pictures of the relatively large, obtained crystals are shown in Fig. 1(b).
Figure 1: Crystal structure of Eu0.8Pt6Al15.8, with the Eu atoms in blue, the Pt atoms in dark gray and the Al atoms in light gray: (a) including polyhedra around the Eu atoms in the $2c$ sites without vacancies (blue) and around the Eu atoms in the $2e$ sites with vacancies (cyan). (b) Photograph of one of the obtained single crystals. (c) Unit cell viewed down the $c$-axis; (d) unit cell viewed down the $a$-axis.
The concentration levels of each element were determined by energy dispersive X-ray spectroscopy (EDS) quantitative chemical analysis using an EDS detector (Thermo NORAN Microanalysis System, model C10001) attached to a JEOL scanning-electron microscope (SEM). An acceleration voltage of 16kV16\ $\text${kV}, working distance of 10mm10\ $\text${mm} and take-off angle of $35^{\circ}$ were used to measure all standards and crystals with unknown composition. Since we were able to obtain polycrystalline arc-melted and single-crystalline samples with the reported composition of Eu2Pt6Al15 [11, 12], flat samples of this phase were used as a standard for Eu, Pt and Al quantification. The spectra were fitted using NIST-DTSA II Microscopium software [17]. The composition of each crystal was measured at four different positions on an arbitrary crystal faces. The average compositions and error bars were obtained from these data, accounting for both inhomogeneity and goodness of fit of each spectra.
[…]
For single-crystal X-ray diffraction (SCXRD) measurements, regularly shaped crystal fragments of Eu0.8Pt6Al15.8 were obtained by fracturing the flux grown crystals. The fragments were glued to quartz fibers using beeswax. The data sets were collected on a Stoe IPDS-II diffractometer (graphite monochromatized Mo $K\alpha$ radiation; oscilation mode). Numerical absorption correction [19] was applied to the data series.
[…]
The temperature-dependent, AC resistance of the samples was measured using a Quantum Design Physical Property Measurement System (PPMS) using the AC transport (ACT) option, with a frequency of 17Hz17\ $\text${Hz} and a 3mA3\ $\text${mA} excitation current. The resistance both parallel and perpendicular to the $c$-axis was measured using a standard four-contact geometry. Electrical contacts with less than $1.5\ \Omega$ resistance were achieved by spot welding $25\ \mu\text{m}$ Pt wire to the samples, followed by adding Epotek H20E silver epoxy, and curing the latter for 1 hour at $120^{\circ}\text{C}$.
3 Results and discussions
3.1 Chemical and structural characterization
Table 1: Single crystal X-ray diffraction measurement and refinement information for Eu0.8Pt6Al15.8.
Chemical formula Eu0.795(7)Pt6Al15.81(6)
Formula Weight 1717.6
Temperature 293 K
Crystal system hexagonal
Space group $P\bar{6}2m$
a=14.2254(6)Åa=14.2254(6)\ $\text{\AA }$
Unit cell dimensions c=8.7363(4)Åc=8.7363(4)\ $\text{\AA }$
Volume 1531.04(12) $\text{\AA }^{3}$
$Z$ 4
Calculated density 7.45 g/cm3
Absorption coefficient 58.7 mm-1
$F(000)$ 2894
[…]
$2\Theta$ range for collection 4.66 to 66.81
Min/Max index [$h$, $k$, $l$] [-21/20, -21/20, -13/13]
Reflections collected 18620
Independent reflections 2204 $R_{int}=0.0543$]
Data/restraint/parameters 2204/0/48
Goodness-of-fit on $F^{2}$ 1.8578
Final $R$ indexes $I\geq 2\sigma(I)$] $R_{1}=0.0304$, $wR_{2}=0.0674$
Final $R$ indexes [all data] $R_{1}=0.0483$, $wR_{2}=0.0675$
Largest diff. peak/hole 14.09 $e^{-}\text{\AA }^{-3}$/-6.15 $e^{-}\text{\AA }^{-3}$
Single-crystal X-ray diffraction measurement on the obtained crystals of Eu0.8Pt6Al15.8, revealed that it adopts a crystal structure Eu1-xPt6Al16-y ($x\sim 0.2$, $y\sim 0.2$) that does not match with any currently reported one. Table 1 shows the information on the measurement and refinement. The system is hexagonal and belongs to the space group $P\bar{6}2m$. A diagram of the unit cell basis of the structure is presented in Fig. 1, and the detailed information of the atomic positions, occupancies and thermal displacement parameters are given in Tab. 2. There are two distinct Eu sites ($2c$ and $2e$) located in the center of the blue and cyan polyhedra defined by the surrounding Pt and Al atoms in Fig. 1(a). The Eu $2e$ sites (in the cyan polyhedra) have an occupancy of around 59(1)%, whereas the other has full occupancy. Besides the Eu1 position, also the Al10 position shows an occupancy lower than 1, while all other positions are fully occupied within three standard deviations. The resulting composition is therefore Eu0.795(7)Pt6Al15.81(6), corresponding to a $x=0.205(7)$ and a $y=0.19(6)$. Since the Eu occupation is close to 1/2, different symmetry reductions were tested, however, no conclusive structural model was obtained in a lower symmetric space group. A potential symmetry loss, however, might be indicative by the remaining residual electron density.
Table 2: Refined atomic positions, site occupancy, and isotropic thermal displacement parameters for Eu0.8Pt6Al15.8, according to SCXRD.
At. Site $x$ $y$ $z$ $U_{\text{iso}}$ Occ.
Eu $2c$ 1/3 2/3 0 0.0093(6) 1
Eu $2e$ 0 0 0.2483(6) 0.008(1) 0.59(1)
Pt $3f$ 0.5262(1) 0 0 0.0062(4) 1
Pt $3f$ 0.8193(2) 0 0 0.0064(6) 1
Pt $3g$ 0.5183(1) 0 1/2 0.0076(4) 1
Pt $3g$ 0.8186(2) 0 1/2 0.0060(6) 1
Pt $12l$ 0.20936(6) 0.45852(6) 0.2483(1) 0.0057(2) 1
Al $3f$ 0.185(2) 0 0 0.011(4) 1
Al $3g$ 0.179(1) 0 1/2 0.008(4) 1
Al $4h$ 1/3 2/3 0.343(2) 0.006(2) 1
Al $6i$ 0.336(1) 0 0.319(2) 0.011(4) 0.87(4)
Al $6j$ 0.1068(8) 0.428(1) 0 0.009(2) 1
Al $6j$ 0.273(1) 0.392(1) 0 0.007(2) 1
Al $6k$ 0.276(1) 0.392(1) 1/2 0.009(2) 1
Al $6k$ 0.163(1) 0.497(1) 0.5 0.017(2) 1
Al $12l$ 0.1159(4) 0.5849(6) 0.243(1) 0.006(1) 1
Al $12l$ 0.1435(6) 0.2571(6) 0.250(1) 0.012(1) 1
The atomic percentages of Eu, Pt and Al were also quantified by EDS and are compared with the atomic percentages given by the SCXRD in Tab. 3. The two results are in agreement within the uncertainties of each analysis, stated in the Tab. 3 as 95% confidence intervals.
Some of the crystals obtained were ground in order to measure powder X-ray diffraction, obtaining the pattern shown in Fig. 2, plotted with black symbols. A Rietveld refinement was performed to fit this pattern, yielding $\chi^{2}=4.5$ and $R_{w}=15.5$. The model structure obtained from the single crystal X-ray refinement was used as the starting point; only the global factor, lattice parameters, atomic positions, microstrains, sample displacement and background were refined. The rest of the parameters, including the thermal displacements and the atomic fractions were kept fixed according to the values reported for the single crystal. The fitted curve is plotted as a red line, showing good agreement with the measured data, as also indicated by the difference (green line).
Figure 2: Powder X-ray diffraction pattern measured for Eu0.8Pt6Al15.8 (black symbols), the best fit obtained by Rietveld refinement (red line) and the residues (green line).
Table 3: Elemental analysis of Eu0.8Pt6Al15.8.
Element from EDS from SCXRD
Eu 3.6(2) 3.51(4)
Pt 28(2) 26.54(7)
Al 68(2) 69.9(8)
3.2 Electronic and magnetic characterization
Fig. 3 shows the ZFC and FC temperature-dependent magnetization normalized by the applied field of 50 Oe, applied along three mutually orthogonal directions: along the $a$-axis (in blue), along the $c$-axis (in green), and along the reciprocal lattice vector $b^{*}$ (in green). These three directions are schematically indicated with arrows on a prism drawing in the main panel. The inset shows an enlarged view of the low temperature range of these results, in order to see the feature more clearly in all three directions of applied field. The magnetization along $a$ and $b^{*}$ shows a downturn upon cooling below 3 K, whereas the magnetization along $c$ shows a more subtle kink and remains higher at low temperatures. The ZFC and FC data remain indistinguishable from each other throughout the measured temperature range.
Figure 3: Main panel: temperature-dependent magnetization of Eu0.8Pt6Al15.8 normalized by the applied field of 50 Oe, oriented along $a$ (blue symbols), $b^{*}$ (red symbols) and $c$ (green symbols). All the shown results were done following both ZFC and FC protocls, which cannot be appreciated in the plot because they coincide with each other. A schematic diagram indicating the different directions with respect to the unit cell was also included. Inset: enlarged view of the low temperature range in order to visualize the feature in greater detail.
The magnetization, in units of Bohr magneton, $\mu_{B}$, per Eu atom is plotted as a function of the applied field along $a$, $b^{*}$ and $c$ in Fig. 4, measured at a constant temperature of 2 K with an increasing magnetic field. As was the case for Fig. 3, the formula unit of Eu0.795(7)Pt6Al15.81(6) was used for computing the intrinsic value of magnetization per Eu. The inset shows an enlarged view of the low-field data, showing a linear behavior that extrapolates to $M(H=0)=0$. No hysteresis is expected in any of the three directions, since no difference was observed between the temperature-dependent ZFC and FC measurements done at 50 Oe (see Fig. 3). As a reference, the black squares in the inset of Fig. 4 show the corresponding ZFC and FC magnetization values extracted from the $M(T)$ measurements at 2 K and a field of 50 Oe applied parallel to $c$, both consistent with the $M(H)$ results. The measurements along $a$ and $b^{*}$ both display some metamagnetic transition around 12 kOe, which is broadened due to the fact that the measurements were taken at a temperature that is close to the transition temperature. Combined, the $M(T)$ and $M(H)$ measurements shown in Figs. 3 and 4 indicate that the sample at low temperatures exhibits antiferromagnetic ordering, without signs of any ferromagnetic component.
The magnetization shown in Fig. 4 does not reach full saturation at the maximum field measured, but it is consistent with a saturation of 7 $\mu_{B}$ expected for Eu2+ ions with $J=7/2$, indicated with a horizontal dashed line in the figure when using the proper stoichiometry of Eu0.795(7)Pt6Al15.81(6). The $\sim 5\%$ difference between the magnetization for $H||a$ (blue curve) and $H||b^{*}$ (red curve) could be due to the background subtraction varying slightly when the Kel-F disk axis is oriented vertically or horizontally  [^1] The background signal was measured independently only for the disk oriented with its axis horizontally (perpendicular to the axis of the straw).
Figure 4: Main panel: Magnetization of Eu0.8Pt6Al15.8 as a function of the applied field, oriented along $a$ (blue), $b^{*}$ (red) and $c$ (green), measured at 2 K. The horizontal dashed line indicates the expected saturation magnetization for Eu2+ ions. Inset: Enlarged view of the low-field results. The vertical dashed line indicates the field at which the $M(T)$ measurements were performed, and the overlapping black squares show the corresponding ZFC and FC magnetization at 2 K obtained from the $M(T)$ measurements done with a field of 50 Oe applied parallel to $c$.
Figure 5 shows the polycrystalline average of the three directions (in black) computed as
$\chi_{\text{ave}}=\frac{M_{a}+M_{b^{*}}+M_{c}}{3H},$ (1)
for $H=50$ Oe which is within the linear regime of the field-dependent magnetization along the three directions. To obtain an accurate value of the transition temperature, $d(\chi_{\text{ave}}T)/dT$ was computed, since this quantity is expected to be proportional to the specific heat for temperatures close to a second-order antiferromagnetic transition [20]. The temperature dependence of $d(\chi_{\text{ave}}T)/dT$ is plotted in magenta in Fig. 5, and the Néel temperature, $T_{N}=2.8(2)$ K, is determined from its maximum, as indicated with the vertical dashed line.
Fig. 6 shows the resistance, normalized to the room temperature value, as a function of temperature, for the current applied along $a$ (blue) and along $c$ (green). The residual resistivity ratio (RRR) is less than 1.5 for both directions, indicating a predominance of temperature-independent defect scattering among other types of scattering, most likely as a consequence of a highly disordered structure. The insets show an enlarged view of the low-temperature range for each direction, and the black arrows indicate the transition temperature according to the magnetic susceptibility measurements. It can be noted that the resistance does not show any discernible feature at the temperature corresponding to the transition in the magnetic susceptibility.
Figure 5: Low temperature polycrystalline average of the magnetic susceptibility of Eu0.8Pt6Al15.8 $\chi_{\text{ave}}$ (black, left axis), and $d(\chi_{\text{ave}}T)/dT$ (magenta, right axis) as a function of temperature.
Figure 6: Main panel: Normalized resistance of Eu0.8Pt6Al15.8 as a function of temperature, measured upon cooling and warming, for the applied current oriented along $a$ (blue) and $c$ (green). Insets: enlarged view of the low temperature range for both measurements; the black arrows indicate the estimated transition temperature according to $d(\chi_{\text{ave}}T)/dT$.
Figure 7: Temperature-dependent magnetization of Eu0.8Pt6Al15.8 normalized by the applied field of 100 Oe applied along the $c$-axis (solid black symbols, left axis), and its inverse (open magenta symbols, right axis). The blue line represent the linear fit applied according to the Curie-Weiss law, in order to obtain the effective magnetic moment, $\mu_{\text{eff}}$, and the Weiss temperature, $\theta_{CW}$.
Magnetic susceptibility was measured up to 300 K applying a field of 100 Oe in order to improve the signal at high temperatures, which decreases asymptotically, and still ensure to lie within the linear regime of $M(H)$, as shown in the inset of Fig. 4 even in the ordered state. It was only measured with the field applied parallel to $c$, since the susceptibility is isotropic in the paramagnetic state, as expected for Eu2+ moments with $L=0$, and as shown in the main panel of Fig. 3. This susceptibility and its inverse are plotted in Fig. 7 in black filled symbols (left axis) and magenta open symbols (right axis), respectively. The latter displays a linear dependence as expected for a Curie-Weiss law behavior. A linear fit was done for $100\leq T\leq 300$ K in order to obtain the effective moment, $\mu_{\text{eff}}$ and the Weiss temperature, $\theta_{CW}$, according to
$\chi^{-1}=\frac{\mu_{\text{eff}}^{2}N_{A}}{3k_{B}}(T-\theta_{CW})$ (2)
This yielded $\mu_{\text{eff}}=8.0(1)\ \mu_{B}$, which agrees with the expected value for Eu2+ ions (7.94 $\mu_{B}$) within uncertainty; and $\theta_{CW}=-5.0(1)$ K, which would indicate the paramagnetic phase has predominantly antiferromagnetic interactions, and is consistent with the low-temperature antiferromagnetic ordering.
4 Conclusion
A ternary compound of Eu, Pt and Al with no reported structural analogs was discovered and characterized. It was obtained in single crystalline form out of a ternary melt with excess Al.
Its chemical formula can be expressed as Eu1-xPt6Al16-x, with $x=0.205(7)$ and $y=0.19(6)$, as determined by refining the occupancy of the single crystal X-ray diffraction data. The chemical composition determined from energy dispersive X-ray spectroscopy was consistent. The structure is hexagonal and belongs to the $P\bar{6}2m$ space group. The characterization measurements done on this phase are consistent with the Eu2+ moments ordering antiferromagnetically below 2.8(2) K.
5 Acknowledgements
J.S. thanks Dolores María Bolo for helping with the graphics design. Work at Ames National Laboratory (J.S., S.L.B., and P.C.C.) was supported by the U.S. Department of Energy, Office of Basic Energy Science, Division of Materials Sciences and Engineering. Ames National Laboratory is operated for the U.S. Department of Energy by Iowa State University under Contract No. DE-AC02-07CH11358. Work at Universitat des Saarlandes (O.J.) was supported by the German Research Foundation DFG (JA 1891-10-1).
6 Data Availability
The data that support the findings of this article are openly available [21]. The .cif file obtained from the single crystal x-ray diffraction measurement and refinement is available as Supplemental Material [22].